3? − 1, −5 ≤ ? < 1 ?(?)= { 4,1≤?≤3 6 − ?, 3 < ? ≤ 5

How do I graph this?
Answer #1

f(x) = 3 -1 -5 << | < x < 3 3<< 5 6 - 2 fux y = 3х-( ч 2 е у = 6-х х 19 - 5 - 16 3 3 1 1 2 a fox) 2 न